A 60kg boy and a 40kg girl decide to play "tug of war" on a frictionless surface. In the process, the girl accelerated toward the boy at 3.0m/sec2. what is the acceleration of the boy toward the girl?
1
Expert's answer
2011-10-18T12:50:41-0400

F1-force which is& appositional by boy
F2-by girl
F1-F2=40kg*3m/s^2
Also F1-F2=60kg*a,
where a is the acceleration of the boy toward the girl
Hence a=40kg*3 m/s^2/60kg =2m/s^2